Find the 100th term of the arithmetic sequence with first term 5 and
eighth term 19
1
Expert's answer
2020-04-13T12:37:46-0400

Nth term of an arithmetic sequence is given by "(a+(n-1)d" )

then "T_1=5" which means a=5

we are given 8th term as 19

"T_8=19\\5+7d=19\\d=2"


We have to find 100th term

"T_{100}=a+(n-1)d=5+99\times2=203"